a) Integrate 1/(1 - x^4) 


b) Integrate 1/(1 + x^4)


 

Grade:11

2 Answers

Varun Acharya
10 years ago
integrate the first one by partial fractions method... (1-x^4) = (1-x^2)(1+x^2) and proceed....
neeraj agarwal
10 years ago
multiply numerator and denominator by 1- x^4 ... put x^4 = sinx